OpenStudy (anonymous):

3x + (4/5) <= 4x + (3/5)

OpenStudy (anonymous):

3x + (4/5) <= 4x + (3/5) 3x + (1/5) <= 4x (1/5) <= x

OpenStudy (vishweshshrimali5):

x>=1/5